HFT Software

2015-01-28 11:00
America/New York
Wednesday, 28 January 2015
EVO, at 16:00 (GMT), duration : 00:00
 Agenda (Xin)

 

1) run 15 status

      - operations  !!! BNL is closed but Bill Christie reminded us that cosmic runs will go on.

      - cosmic run statistics, calibration status, alignment production:

[Regarding the cosmic run statistics, as of this morning, we have taken 
-Zero Field data,  PXL+IST+SSD+TPC,  94 hours   (there are additional 22 hours of data without TPC)

 -Full Field data,  PXL+IST+SSD+TPC,  58 hours]

 

2) run 14 production readiness

      - HFT calibration - status table verification request sent out
https://drupal.star.bnl.gov/STAR/blog/dongx/production-request-run14-pxlist-calibration-table-verifications

      - tracking efficiency study update - Hao/Michael/Anju

      - other calibration status - TPC etc.

 

3) embedding meeting last Thursday - progress?

 

4) others?

TimeTalkPresenter
11:00Minutes ( 00:20 ) 0 files